lim (x2-4)/x-2

s→2




1
Expert's answer
2022-04-05T14:20:22-0400
"\begin{aligned}
\lim_{x\rightarrow2}\frac{x^2-4}{x-2}=&
\lim_{x\rightarrow2}\frac{(x-2)(x+2)}{x-2}\\
=&\lim_{x\rightarrow2}({x+2})\\
=&2+2\\
\lim_{x\rightarrow2}\frac{x^2-4}{x-2}= &4
\end{aligned}"